OpenStudy (anonymous):

lnx - ln5 =4 how would I solve show process please

OpenStudy (aravindg):

Use \[\log a -\log b=\log \dfrac{a}{b}\]

OpenStudy (anonymous):

umm how would i go about this?

OpenStudy (aravindg):

Just compare the terms with the equation I gave above lnx - ln5 =?

OpenStudy (anonymous):

ln -x/5?

OpenStudy (anonymous):

im kinda confused what to do with the 4

OpenStudy (aravindg):

log x/5=4 This means e^4=x/5 You can solve for x now!

OpenStudy (anonymous):

ohh I understand now so x=5e^4

OpenStudy (solomonzelman):

AravindG, I thought it would become\[\ln(\frac{x}{5})=4\](not log, but ln, with base e)

OpenStudy (aravindg):

Yes I was referring to natural log. You can see that I took e in the end though.So no worries :)

OpenStudy (aravindg):

Yes @itsJonLe :)

OpenStudy (solomonzelman):

Yeah :)

OpenStudy (anonymous):

okay thank you very much!! :DD

OpenStudy (aravindg):

yw :)